answer the question below about the quadratic function g(x) =2x^2+8x+12
1) does the function have a minium or maxium value
2) what is the funtion mininum or maxium value
3) where does the minimum or maxium occur

Answer by nerdybill(7384)

answer the question below about the quadratic function g(x) =2x^2+8x+12
1) does the function have a minium or maxium value
Because the leading coefficient is POSITIVE, the parabola opens downwards -- thus the function has a maximum.
.
It is easier to answer part 3) first:
3) where does the minimum or maxium occur
this is the "axis of symmetry":
x = -b/(2a)
x = -8/(2*2)
x = -8/(4)
x = -2
.
2) what is the funtion mininum or maxium value
plug -2 into the the equation to find the max:
g(x) =2x^2+8x+12
g(-2) =2(-2)^2+8(-2)+12
g(-2) =2(4)+(-16)+12
g(-2) =8-16+12
g(-2) =-8+12
g(-2) = 4